if A=(1 3 2 1),then find the determinant of A 2 -2a is

if A=(1 3


       2 1),then find the determinant of A2 -2a is

Grade:

1 Answers

ramya ramanathan
11 years ago

A^2 = (7      6

           4      7)

-2A = (-2 -6

          -4 -2)

A^2-2A = (5 0

                0 5)

|A^2-2A| = 25

 

I am new to askiitians.com. Sorry if the answer is wrong.